### v3

@@ 1.1 Bio-inspired autonomous navigation
While these projects design systems whose control is inspired from honeybee behavior, our work uses a learning algorithm to train agents with sensory input similar to a bee to fly. Our goal is to inspect the learned behavior to determine what the agent is paying most attention to in its sensory input. Since there is evidence supporting the theory that bees use optic flow from different areas of vision for different control tasks [24, 25, 26, 27], we investigate whether similar patterns arise from learning agents that use the same sensory information.
@@ 3.1 Observation Space
Explicitly, an element of the observation space $\Omega$ is an array containing the optic flow magnitude and orientation at each pixel in the camera’s field of view. We represent this as an ‘image’ with three channels: the optic flow magnitude, and the horizontal and vertical components of the normalized optic flow vector. We set angular velocity to zero during optic flow calculation ($W=\mathbf{0}$), since we found that the contribution from the drone’s angular velocity is often much larger than the contribution from its linear motion. [^3] This can be simulated on physical systems either through measuring angular velocity, or through estimating it from observed optic flow. We visualize an observation in Fig. LABEL:fig:pay-attention(b), where the agent is traveling forward and right. Here, we display (unnormalized) optic flow vectors as black arrows over a heatmap of the optic flow magnitude.
@@ 4 Results
We train four agents independently in this way, so that we may inspect general trends in attention patterns rather than the learned attention of any one specific trained agent. We sample a few trajectories from one of these trained agents navigating a tunnel. We use the recorded optic flows at each timestep in these trajectories to obtain SHAP values for each of the four agents on these observations. We take the absolute value of these SHAP values, apply Gaussian smoothing, and take the average across all four agents. The resulting attention patterns for a single timestep are displayed in Figures LABEL:fig:pay-attention and 5.
We find that the trained agents are successful in navigating the tunnels using optic flow observations. Their trajectories (displayed in Fig. LABEL:fig:bee-flying) appear to center the agent in the non-obstructed region of the tunnel.
We find that regions of discontinuity in optic flow magnitude make a disproportionally larger contribution to the agent’s decisions. This is seen in Fig. LABEL:fig:pay-attention, where the agents pay most attention to the nearest edges of an obstacle. Aside from this, the agents seem to pay attention to regions of high optic flow, such as the center of the nearest obstacle and the closest parts of the tunnel walls in Fig. LABEL:fig:pay-attention. We observe this attention pattern throughout the trajectories, in addition to the example shown.
@@ 5 Discussion
In our experiments, our trained agents appear to balance the optic flow perceived in the left and right regions of their visual field to maintain a centered flight path. We find that the agents pay most attention to nearby edges of obstacles, where discontinuities in optic flow are most significant (Fig. 5). This results in the the well documented centering behavior of honeybees [62, 63] appearing in the agents’ trajectories (Fig. LABEL:fig:bee-flying), as well as detection and obstacle avoidance evident from lateral deviation and reduction in forward speed. The similarities of the trajectories to navigation patterns found in bees [12, 6] indicate that real bees also respond strongly to discontinuities in optic flow, such as those induced by the edges of obstacles. This also suggests that an effective computational navigation strategy should involve discontinuity detection, as opposed to regulating raw optic flow values.

### v4

@@ 1.1 Bio-inspired autonomous navigation
While these works design systems whose control is inspired from honeybee behavior, our work uses a learning algorithm to train agents with sensory input similar to a bee to fly. Our goal is to inspect the learned behavior to determine what the agent is paying most attention to in its sensory input. Since there is evidence supporting the theory that bees use optic flow from different areas of vision for different control tasks [36, 38, 24, 51], we investigate whether similar patterns arise from learning agents that use the same sensory information.
@@ 3.1 Observation Space
Explicitly, an element of the observation space $\Omega$ is an array containing the optic flow magnitude and orientation at each pixel in the camera’s field of view. We represent this as an ‘image’ with three channels: the optic flow magnitude, and the horizontal and vertical components of the normalized optic flow vector. We set angular velocity to zero during optic flow calculation ($W=\mathbf{0}$), since we found that the contribution from the drone’s angular velocity is often much larger than the contribution from its linear motion. [^3] This can be simulated on physical systems either through measuring angular velocity, or through estimating it from observed optic flow. We visualize an observation in Fig. 5(a)(b), where the agent is traveling forward and right. Here, we display (unnormalized) optic flow vectors as black arrows over a heatmap of the optic flow magnitude.
@@ 4 Results
We train four agents independently in this way, so that we may inspect general trends in attention patterns rather than the learned attention of any one specific trained agent. We sample a few trajectories from one of these trained agents navigating a tunnel. We use the recorded optic flows at each timestep in these trajectories to obtain SHAP values for each of the four agents on these observations. We take the absolute value of these SHAP values, apply Gaussian smoothing, and take the average across all four agents. The resulting attention patterns for a single timestep are displayed in Figures 5(a) and 5.
(a) The attention pattern of trained agents flying through a tunnel, averaged over four independently trained agents.
We find that the trained agents are successful in navigating the tunnels using optic flow observations. Their trajectories (displayed in Fig. 5(b)) appear to center the agent in the non-obstructed region of the tunnel.
(b) Trajectories from a trained agent flying through different tunnels: (a) Easier tunnel with two obstacles placed off-center, (b) Difficult tunnel with three obstacles, one placed centrally and two placed off-center.
We find that regions of discontinuity in optic flow magnitude make a disproportionally larger contribution to the agent’s decisions. This is seen in Fig. 5(a), where the agents pay most attention to the nearest edges of an obstacle. Aside from this, the agents seem to pay attention to regions of high optic flow, such as the center of the nearest obstacle and the closest parts of the tunnel walls in Fig. 5(a). We observe this attention pattern throughout the trajectories, in addition to the example shown.
@@ 5 Discussion
In our experiments, our trained agents appear to balance the optic flow perceived in the left and right regions of their visual field to maintain a centered flight path. We find that the agents pay most attention to nearby edges of obstacles, where discontinuities in optic flow are most significant (Fig. 5). This results in the the well documented centering behavior of honeybees [53, 3] appearing in the agents’ trajectories (Fig. 5(b)), as well as detection and obstacle avoidance evident from lateral deviation and reduction in forward speed. The similarities of the trajectories to navigation patterns found in bees [54, 58] indicate that real bees also respond strongly to discontinuities in optic flow, such as those induced by the edges of obstacles. This also suggests that an effective computational navigation strategy should involve discontinuity detection, as opposed to regulating raw optic flow values.